OpenStudy (anonymous):

i have the same question that kailynn has but what number do i multiply by 100 to get the percentage?

OpenStudy (texaschic101):

whats the question ? Can you please post it

OpenStudy (anonymous):

A bag contains 30 red tiles, 15 green tiles, and 5 yellow tiles. One tile is drawn and then replaced. Then a second tile is drawn. What is the probability that the first tile is yellow and the second tile is green? A. 1% B. 3% C. 6% D. 18%

OpenStudy (texaschic101):

I am not good at probabilities....I know you add all the marbles. 30 + 15 + 5 = 50 first tile being yellow is 5/50 or 1/10 second tile being green is 15/50 or 3/10 1/10 * 3/10 = 3/100 = .03 = 3% you might need to get a second opinion on this...this is mainly a guess